### v1

@@ 2.2.1 The architecture of functional autoencoder
where pool is a pooling operation invariant to the order of its inputs—in our case, the average pooling. The average pooling operation reduces the discrepancy between latent representations derived from sparse and dense inputs, thereby endowing the functional autoencoder with the capability for accurate sparse-data reconstruction.
@@ 2.2.2 Training objective and masked training
[…]
2.3 Conditional diffusion model for detail generation
After reconstructing the principal structures with the functional autoencoder, a conditional diffusion model is trained to generate the small-scale components, defined as the difference between the ground truth and the reconstructed principal structures, conditioned on the principal structures. By defining the detail or residual $\boldsymbol{d}=\boldsymbol{u}-\boldsymbol{m}$, learning $p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})$ is equivalent to learning $p(\boldsymbol{d}|\boldsymbol{m},\boldsymbol{y})$ as it is a deterministic relationship. Modeling the residual $\boldsymbol{d}$ simplifies the learning task, reduces complexity of modeling, and improves training stability and convergence. It also introduces an inductive bias, constraining the reconstructed $\boldsymbol{u}$ to remain close to $\boldsymbol{m}$, thereby restricting the solution space and mitigating ill-posedness. As discussed above, the functional autoencoder approximates $p(\boldsymbol{m}|\boldsymbol{y})$ with a Dirac delta function $\delta(\boldsymbol{m}-\hat{\boldsymbol{m}}(\boldsymbol{y}))$. Consequently, the conditional diffusion model should be trained to model:
$p(\boldsymbol{u}|\boldsymbol{y})=\int p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})\,p(\boldsymbol{m}|\boldsymbol{y})\,d\boldsymbol{m}$ $=\int p(\boldsymbol{d}|\boldsymbol{m},\boldsymbol{y})\,p(\boldsymbol{m}|\boldsymbol{y})\,d\boldsymbol{m}\approx\int p(\boldsymbol{d}|\boldsymbol{m},\boldsymbol{y})\,\delta(\boldsymbol{m}-\hat{\boldsymbol{m}}(\boldsymbol{y}))\,d\boldsymbol{m}=p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y}).$
@@ 2.3.1 Conditional denoising diffusion probabilistic model
Training of denoising diffusion probabilistic models (DDPMs) consists of a diffusion process and a denoising process. Diffusion process transforms $\boldsymbol{d}_{0}$ from the real data distribution $\boldsymbol{d}_{0}\sim p(\boldsymbol{d}_{0})$ into a pure Gaussian noise $\boldsymbol{d}_{t}\sim\mathcal{N}(0,\boldsymbol{I})$ by successively applying the following Markov diffusion kernel:
@@
Given $\boldsymbol{d}_{t}$, the denoising process aims to recover $\boldsymbol{d}_{0}$ by recursively learning the transition from $\boldsymbol{d}_{t-1}$ to $\boldsymbol{d}_{t}$, which is defined as the following Gaussian distribution:
@@
However, this unconditional scheme fails to preserve the uniqueness of reconstruction due to the lack of guidance during generation. Therefore, $\hat{\boldsymbol{m}}(\boldsymbol{y})$ is injected as the condition for the denoising process:
@@
2.3.2 Manifold constrained gradient
The conditional distribution $p(\boldsymbol{d}_{t-1}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}))$ has been modeled by training the conditional diffusion model. The $p(\boldsymbol{d}_{t-1}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$ should be modeled to further preserve the uniqueness of reconstruction and impose measurement consistency. However, directly model the condtional distribution $p(\boldsymbol{d}_{t-1}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$ is impossible due to the immense variability of sparse measurement patterns. We model it during inference instead of training of conditional diffusion model. Specifically, the Manifold Constrained Gradient (MCG) [chung2022improving] is used to inject an additional gradient term to guide the generation towards observed measurements without retraining the model. This gradient refers to the score function in score-based generative models (SGMs) [song2020score], defined as the gradient of the log-density of the data distribution $\nabla_{\boldsymbol{d}_{t}}\log p_{t}(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}))$.
Although DDPMs and SGMs were originally proposed under different formulations—predicting added noise in DDPMs and estimating score functions in SGMs—it has been shown that they are theoretically equivalent under certain parameterizations [song2020score]. Specifically, when a DDPM is trained to predict the Gaussian noise $\boldsymbol{\epsilon}_{\theta}$ added at each step, the network implicitly learns the score function of the perturbed data distribution $p_{t}(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}))$ up to a scaling factor:
@@ 2.3.2 Manifold constrained gradient
Therefore, the sampling equation in Eq. 10 can be reformulated as:
$\mathbf{\boldsymbol{d}}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{\boldsymbol{d}}_{t}+(1-\alpha_{t})\nabla_{\boldsymbol{d}_{t}}\log p_{t}(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}))\right)+\sigma_{t}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (12)
@@
where $\nabla_{\boldsymbol{d}_{t}}\log p_{t}(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}))$ has been modeled by conditional diffusion model. To approximate the likelihood term $p_{t}(\boldsymbol{y}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}))$, we assume that it follows a Gaussian form, i.e.,
$p_{t}(\boldsymbol{y}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}))\propto\exp\left(-\frac{1}{2\sigma_{c}^{2}}\|\boldsymbol{y}-\mathcal{M}(\hat{\boldsymbol{m}}(\boldsymbol{y})+\boldsymbol{d}_{t})\|_{2}^{2}\right),$ (14)
where $\mathcal{M}(\cdot)$ denotes the mask operator and $\sigma_{c}^{2}$ denotes the variance associated with measurement noise. $\sigma_{c}^{2}$ is set as 10000 in all experiments of this study. Under this assumption, its gradient with respect to $\boldsymbol{d}_{t}$ yields:
@@
Therefore, the sampling formula without retraining the model is given as:
[…]
However, it drives the intermediate states at each denoising step to match the observations, which pulls the sample path away from the learned data manifold. Chung et al. [chung2022improving] reveal that the Bayes optimal denoising step from the Tweedie’s formula [robbins1992empirical] leads to a preferred condition for diffusion model both empirically and theoretically. As a result, Tweedie’s formula is introduced here to impose the measurement consistency while improving the generation accuracy. In the case of Gaussian noise, a classic result of Tweedie’s formula tells us that one can achieve the denoised result by computing the posterior expectation:
@@
2.3.3 Mask-cascade training
We propose a mask-cascade training strategy to improve the robustness of the conditional diffusion model under varying sparsity patterns. A key challenge is that the functional autoencoder, when provided with different sparse samplings of the same field, reconstructs slightly different principal structures. Since it is infeasible to enumerate all possible reconstructions for training, directly using a fixed dataset would restrict the generalization ability of the diffusion model.
To address this issue, our strategy explicitly integrates the stochasticity of sparse sampling into training the conditional diffusion model. As illustrated in Fig. 1(b), once the functional autoencoder is fully trained and its parameters are frozen, the training of the conditional diffusion model proceeds as follows: in each training step, we randomly generate a sparse mask and apply it to the ground-truth field $\boldsymbol{u}$ to obtain a partial observation $\boldsymbol{y}$. This observation is then passed through the functional autoencoder to reconstruct an approximate principal structure $\hat{\boldsymbol{m}}(\boldsymbol{y})$. The conditional diffusion model is then conditioned on $\hat{\boldsymbol{m}}(\boldsymbol{y})$, and trained to generate the fine-scale component $\boldsymbol{d}=\boldsymbol{u}-\hat{\boldsymbol{m}}(\boldsymbol{y})$. By repeatedly sampling different masks across training steps, the conditional diffusion model is continuously exposed to a diverse ensemble of approximate principal structure reconstructions. This effectively augments the training distribution of conditioning inputs without the need for explicitly precomputing or storing all possible variants. As a result, the conditional diffusion model learns to model the conditional distribution $p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}))$, thereby achieving improved robustness and consistency across varying sparsity patterns.
As shown in Fig.1(c), the inference process of autoencoder-diffusion cascade proceeds as follows. Given arbitrary sparse observations $\boldsymbol{y}$, the functional autoencoder first reconstructs the principal structure $\hat{\boldsymbol{m}}(\boldsymbol{y})$. This reconstructed structure is then provided as a condition to the conditional diffusion model, which generates the fine-scale details $\hat{\boldsymbol{d}}$ through the manifold constrained gradient-enhanced sampling scheme in Eq. (18). Finally, the full-field reconstruction is obtained by combining the two components $\hat{\boldsymbol{u}}=\hat{\boldsymbol{m}}(\boldsymbol{y})+\hat{\boldsymbol{d}}$. It is worth noting that, during inference, Cas-Sensing can still achieve reliable reconstruction even when the sampling rate of sparse observations differs significantly from the rates used in mask-cascade training.
(a) The masked training scheme of functional autoencoder. The trained functional autoencoder is to approximate $p(\boldsymbol{m}|\boldsymbol{y})$ with $\delta(\boldsymbol{m}-\hat{\boldsymbol{m}}(\boldsymbol{y}))$ as it learns the deterministic mapping between $\boldsymbol{y}$ and $\boldsymbol{m}$. $x_{e}$ and $u_{e}$ denote the coordinates and field values fed into the encoder, while $x_{d}$ and $u_{d}$ represent the coordinates and reconstructed values produced by the decoder.
(b) The mask-cascade training scheme of conditional diffusion model, which ensure the conditional diffusion model’s accurate capturing $p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}))$.
(c) The inference scheme of Cas-Sensing based on an autoencoder-diffsusion cascade, which is the process of sampling from $p(\boldsymbol{u}|\boldsymbol{y})\approx p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$. Manifold Constrained Gradient (MCG) is used to impose measurement consistency and sample from $p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$ without retraining the model.
Figure 1: Overview of the proposed Cascaded Sensing (Cas-Sensing) based on an autoencoder-diffusion cascade.
@@ 3.1 Reconstructing cylinder flow velocity fields with simulation data
To evaluate the effectiveness of the proposed Cas-Sensing based on an autoencoder-diffusion cascade in reconstructing full-field data from randomly sparse observations under varying geometric boundary conditions, we first constructed a 2D cylinder flow dataset. This dataset captures incompressible fluid dynamics around a 2D circular cylinder within a channel:
@@
Specifically, the computational domain is defined as a rectangular region of size $1.6\times 0.6$, within which both the cylinder center and radius are randomly sampled. This procedure yields a dataset comprising 96 distinct boundary configurations, each associated with 100 temporal snapshots. All flow fields are interpolated onto a uniform grid of size $256\times 96$. The visualization of a sample from the Cylinder Flow is presented in Fig. 2. For the purpose of model training and evaluation, the dataset is partitioned based on boundary configurations rather than temporal snapshots, ensuring that the test set contains unseen geometrical variations. Specifically, 76 boundary configurations, corresponding to a total of 7600 flow field snapshots, are employed as the training set, while the remaining 20 configurations, consisting of 2000 snapshots, are reserved exclusively for testing.
@@
We first employ a functional autoencoder to learn compact representations of the circular cylinder flow fields. To encourage robustness, the model is trained in a masked manner, where only $50\%$ of the grid points in each snapshot are provided to the encoder and the task is to reconstruct the full field from this partial information. To investigate the latent representation learned by the functional autoencoder, we first visualize the extracted latent vectors of the cylinder flow dataset. Specifically, the 32-dimensional latent vectors are projected into a two-dimensional space using the t-distributed Stochastic Neighbor Embedding (t-SNE) technique [maaten2008visualizing]. Each point in Fig. 3 corresponds to the latent variable of a single flow snapshot. It can be observed that the training data associated with different internal boundary configurations are well separated in the latent space, and the latent vectors of snapshots from the same configuration form circular patterns. This circular structure reflects the inherent periodicity of the velocity fields in cylinder wake flows. Moreover, the latent vectors corresponding to the test dataset exhibit the same characteristic structures, indicating that the functional autoencoder has effectively captured both the temporal and geometric features of the flow fields.
@@
Figure 4: The reconstructed principal structures and kernel density estimates for RMSE on the reference sample across 100 randomly chosen meshes.

### v2

@@ 2.2.1 The architecture of functional autoencoder
where pool denotes a pooling operation that is invariant to the order of its inputs—specifically, average pooling in our implementation. Owing to its permutation-invariant property, average pooling enables the functional autoencoder to process arbitrarily ordered and distributed sparse inputs. Moreover, it can adaptively attenuating high-frequency variations while preserving low-frequency components, which is consistent with the well-known smoothing effect of averaging operations. Since average pooling reduces the discrepancy between latent representations obtained from sparse and dense observations, the learned latent space becomes largely insensitive to the input sampling ratio, endowing the functional autoencoder with robust and accurate reconstruction capability under extreme sparsity.
@@ 2.2.2 Training objective and masked training
where the second term, $\mathcal{E}(u;\zeta)||_{2}^{2}$ with coefficient $\beta$, plays the role of regularisation, ensuring that the latent representation is smooth and structured.
@@
2.3 Conditional diffusion models for refined-scale detail generation
In Cas-Sensing, the second stage is designed to generate refined-scale details conditioned on the recovered coarse-scale structures, while explicitly enforcing consistency with sparse measurements. To this end, we adopt a Denoising Diffusion Probabilistic Model (DDPM) as the generative backbone, owing to its Bayesian interpretation of the sampling process, which enables additional conditioning constraints to be injected at inference time without retraining the model. Specifically, the reverse diffusion process can be viewed as sampling from a data-driven prior, and measurement consistency can be incorporated by reformulating unconditional sampling as posterior sampling under an observation likelihood. This property allows sparse observational constraints to be enforced through gradient-based corrections during denoising (e.g., manifold-constrained gradients), achieving data fidelity while preserving the learned data manifold, and making DDPMs particularly suitable for deployment under varying sensor configurations and sparsity levels.
After reconstructing the coarse-scale structures $\boldsymbol{m}$ with the functional autoencoder, the conditional diffusion model is trained to generate the refined-scale components defined as the residual $\boldsymbol{d}=\boldsymbol{u}-\boldsymbol{m}$. Under this deterministic decomposition, learning the conditional distribution $p(\boldsymbol{u}\mid\boldsymbol{m},\boldsymbol{y})$ is equivalent to learning $p(\boldsymbol{d}\mid\boldsymbol{m},\boldsymbol{y})$. Modeling the residual significantly simplifies the learning task by reducing distributional complexity, improving training stability and convergence, and introducing an inductive bias that constrains the reconstructed field to remain close to the coarse-scale estimate. As discussed in Eq. (2), the functional autoencoder approximates $p(\boldsymbol{m}\mid\boldsymbol{y})$ with a Dirac delta distribution $\delta(\boldsymbol{m}-\hat{\boldsymbol{m}}(\boldsymbol{y}))$, leading to the following cascaded probabilistic formulation:
$p(\boldsymbol{u}\mid\boldsymbol{y})=\int p(\boldsymbol{u}\mid\boldsymbol{m},\boldsymbol{y})\,p(\boldsymbol{m}\mid\boldsymbol{y})\,d\boldsymbol{m}\approx\int p(\boldsymbol{d}\mid\boldsymbol{m},\boldsymbol{y})\,\delta(\boldsymbol{m}-\hat{\boldsymbol{m}}(\boldsymbol{y}))\,d\boldsymbol{m}=p(\boldsymbol{d}\mid\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y}).$ (7)
@@ 2.3.1 Conditional denoising diffusion probabilistic model
Training of DDPMs consists of a diffusion process and a denoising process. Diffusion process transforms $\boldsymbol{d}_{0}$ from the real data distribution $\boldsymbol{d}_{0}\sim p(\boldsymbol{d}_{0})$ into a pure Gaussian noise $\boldsymbol{d}_{t}\sim\mathcal{N}(0,\boldsymbol{I})$ by successively applying the following Markov diffusion kernel:
@@
Reversely, given $\boldsymbol{d}_{t}$, the denoising process aims to recover $\boldsymbol{d}_{0}$ by recursively learning the transition from $\boldsymbol{d}_{t-1}$ to $\boldsymbol{d}_{t}$, which is defined as the following Gaussian distribution:
@@
However, this unconditional scheme fails to preserve the uniqueness of reconstruction due to the lack of guidance during generation. Therefore, the coarse-scale inference $\hat{\boldsymbol{m}}(\boldsymbol{y})$ is injected as the condition for the denoising process:
@@
2.3.2 Training stage: Mask-cascade training
Figure 3: The proposed mask-cascade training (MCT) of the conditional diffusion model.
A key challenge is that the functional autoencoder, when provided with different sparse samplings from the same field, reconstructs slightly different coarse-scale structures $\hat{\boldsymbol{m}}(\boldsymbol{y})$. Since it is infeasible to enumerate all possible reconstructions for training, directly using a fixed dataset would restrict the generalization ability of the diffusion model. In this regard, we propose a mask-cascade training strategy to improve the robustness of the conditional diffusion model under varying sparsity patterns.
Our strategy explicitly integrates the stochasticity of sparse sampling into training the conditional diffusion model. As illustrated in Fig. 3, once the functional autoencoder is fully trained and its parameters are frozen, the training of the conditional diffusion model proceeds as follows: in each training step, we randomly generate a sparse mask with fixed mask ratio and apply it to the ground-truth field $\boldsymbol{u}$ to obtain a partial observation $\boldsymbol{y}$. Notably, the mask ratio here is chosen to be significantly lower than that used for masked training the functional autoencoder, so as to explicitly train the conditional diffusion model to operate under extremely sparse observations. This observation is then passed through the trained (fixed) functional autoencoder to reconstruct an approximate coarse-scale structure $\hat{\boldsymbol{m}}(\boldsymbol{y})$. The conditional diffusion model is then conditioned on $\hat{\boldsymbol{m}}(\boldsymbol{y})$, and trained to generate the fine-scale component $\boldsymbol{d}=\boldsymbol{u}-\hat{\boldsymbol{m}}(\boldsymbol{y})$. By repeatedly implementing different masks across training steps, the conditional diffusion model is continuously exposed to a diverse ensemble of approximate coarse-scale structure reconstructions. This effectively augments the training distribution of conditioning inputs without the need for explicitly precomputing or storing all possible variants. As a result, the conditional diffusion model learns to model the conditional distribution $p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}))$, thereby achieving improved robustness and consistency across varying sparsity patterns.
2.3.3 Inference stage: Manifold constrained gradient
The conditional distribution $p(\boldsymbol{d}_{t-1}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}))$ has been modeled by training the conditional diffusion model. According to the formulation in Eq.(7), the $p(\boldsymbol{d}_{t-1}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$ should be modeled to further preserve the uniqueness of reconstruction and impose measurement consistency. However, directly model the condtional distribution $p(\boldsymbol{d}_{t-1}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$ is impossible due to the immense variability of sparse measurement patterns. We model it during the inference stage, instead of training of such a conditional diffusion model. Specifically, the Manifold Constrained Gradient (MCG) [chung2022improving] is used to inject an additional gradient term to guide the generation towards observed measurements without retraining the model. This gradient refers to the score function in score-based generative models (SGMs) [song2020score], defined as the gradient of the log-density of the data distribution $\nabla_{\boldsymbol{d}_{t}}\log p_{t}(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}))$.
Although DDPMs and SGMs were originally proposed under different formulations – predicting added noise in DDPMs and estimating score functions in SGMs – it has been shown that they are theoretically equivalent under certain parameterizations [song2020score]. Specifically, when a DDPM is trained to predict the Gaussian noise $\boldsymbol{\epsilon}_{\theta}$ added at each step, the network implicitly learns the score function of the perturbed data distribution $p_{t}(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}))$ up to a scaling factor:
@@ 2.3.3 Inference stage: Manifold constrained gradient
where $\nabla_{\boldsymbol{d}_{t}}\log p_{t}(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}))$ has been modeled by conditional diffusion model. To approximate the likelihood term $p_{t}(\boldsymbol{y}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}))$, we assume that it follows a Gaussian form. Under this assumption, its gradient with respect to $\boldsymbol{d}_{t}$ yields:
@@
where $\mathcal{M}(\cdot)$ denotes the mask operator and $\sigma_{c}^{2}$ denotes the variance associated with measurement noise. $\sigma_{c}^{2}$ is set as 10000 in all experiments of this study.
According to the Bayes rule in Eq. 14, the measurement consistency can be injected into the trained diffusion model by adding this new gradient into the sampling formulation in Eq. 12. However, it drives the intermediate states at each denoising step to match the observations, which pulls the sample path away from the learned data manifold. Chung et al. [chung2022improving] reveal that the Bayes optimal denoising step from the Tweedie’s formula [robbins1992empirical] leads to a preferred condition for diffusion model both empirically and theoretically. As a result, Tweedie’s formula is introduced here to impose the measurement consistency while improving the generation accuracy. In the case of Gaussian noise, a classic result of Tweedie’s formula tells us that one can achieve the denoised result by computing the posterior expectation:
@@
As shown in Fig.4, the inference process of autoencoder-diffusion cascade proceeds as follows. Given arbitrary sparse observations $\boldsymbol{y}$, the functional autoencoder first reconstructs the coarse-scale structure $\hat{\boldsymbol{m}}(\boldsymbol{y})$. This reconstructed structure is then provided as a condition to the conditional diffusion model, which generates the fine-scale details $\hat{\boldsymbol{d}}$ through the manifold constrained gradient-enhanced sampling scheme in Eq. (18). Finally, the full-field reconstruction is obtained by combining the two components $\hat{\boldsymbol{u}}=\hat{\boldsymbol{m}}(\boldsymbol{y})+\hat{\boldsymbol{d}}$. It is worth noting that, during inference, Cas-Sensing can still achieve reliable reconstruction even when the sampling rate of sparse observations differs significantly from the rates used in mask-cascade training.
Figure 4: The inference scheme of Cascaded Sensing. Given different sparse measurements of the same target physical field, the framework samples from the conditional distribution $p(\boldsymbol{u}|\boldsymbol{y})\approx p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$, producing multiple reconstruction samples. Measurement consistency is imposed with Manifold Constrained Gradient (MCG) during diffusion sampling, without retraining the model.
@@ 3.1 Reconstructing cylinder flow velocity fields with simulation data
To evaluate the effectiveness of the proposed Cas-Sensing in reconstructing full-field data from randomly sparse observations under varying geometric boundary conditions, we first constructed a synthetic 2D cylinder flow dataset. This dataset captures incompressible fluid dynamics around a 2D circular cylinder within a channel:
@@
Specifically, the computational domain is defined as a rectangular region of size $1.6\times 0.6\ m$, within which both the cylinder center position and radius value are randomly sampled. This procedure yields a dataset comprising 96 distinct boundary configurations, each associated with 100 temporal snapshots. All flow fields are interpolated onto a uniform grid of size $256\times 96$. The visualization of a sample from the Cylinder Flow is presented in Fig. 5.
For the purpose of model training and evaluation, the dataset is partitioned based on boundary configurations rather than temporal snapshots, ensuring that the test set contains unseen geometrical variations. Specifically, 76 boundary configurations, corresponding to a total of 7600 flow field snapshots, are employed as the training set, while the remaining 20 configurations, consisting of 2000 snapshots, are reserved exclusively for testing.
@@
We first employ a functional autoencoder to learn compact representations of the circular cylinder flow fields. To encourage robustness, the model is trained in a masked manner, where only $50\%$ of the grid points in each snapshot are provided to the encoder and the task is to reconstruct the full field from this partial information.
To investigate the latent representation learned by the functional autoencoder, we first visualize the extracted latent vectors of the cylinder flow dataset. Specifically, the 32-dimensional latent vectors are projected into a two-dimensional space using the t-distributed Stochastic Neighbor Embedding (t-SNE) technique [maaten2008visualizing]. Each scattered point in Fig. 6 corresponds to the latent variable of a single flow snapshot. It can be observed that the training data associated with different internal boundary configurations are well separated in the latent space, while the latent vectors of snapshots from the same configuration form circular patterns. This circular structure reflects the inherent periodicity of the velocity fields in cylinder wake flows. Moreover, the latent vectors corresponding to the test dataset exhibit the same characteristic structures, indicating that the functional autoencoder has effectively captured both the temporal and geometric features of the given flow fields.
@@
Figure 7: The reconstructed coarse-scale structures and kernel density estimates for RMSE on the reference sample across 100 randomly chosen meshes.